# Time-reversed information flow through a wormhole in scalar-tensor gravity

### v1

@@ Abstract
This Letter aims to advance new properties of a class of Closed Timelike Curves recently discovered in scalar-tensor gravity [Universe 9, 467 (2023)]. Therein, it was shown that a wormhole acts a gateway between two time-mirrored worlds, where the two asymptotically flat sheets in the Kruskal-Szekeres diagram are glued antipodally along three directions—time $t$ and the polar and azimuth angles $(\theta,\,\varphi)$ of the 2-sphere—to form a wormhole throat. This contrasts with the standard embedding diagram which typically glues the sheets only along the $\theta$ and $\varphi$ directions. Crucially, due to the ‘gluing’ along the $t$ direction, the wormhole becomes a portal connecting the two spacetime sheets with opposite physical time flows, enabling the emergence of closed timelike loops which straddle the throat. We shall point out that this portal mathematically permits the possibility of backward propagation of information against time. This feature is ubiquitous for wormholes in scalar-tensor theories. In addition, we formulate the Feynman sum for transition amplitudes of microscopic particles in the proximity of a wormhole throat in which we account for timelike paths that experience time reversal.
@@
Introduction—Closed timelike curves (CTCs) are mathematical constructs first emerging in the context of General Relativity via the pioneering works of van Stockum and Gödel [1, 2]. These curves represent paths through spacetime that loop back on themselves, allowing an object or observer to return to an event in its own past. Interests in CTCs were revived in the works of Thorne and Novikov in the context of the Morris-Thorne wormholes which are thought to occur in the presence of exotic matter (or in modified gravity) [3, 4, 6, 5, 7, 8]. Despite the problematic nature of CTCs, the mathematical validity sparked considerable interest among gravitational theorists [9, 16, 12, 11, 13, 14, 10, 15, 17, 18, 19]. The concept of CTCs serves as a “gedanken experiment”, compelling us to grapple with and reevaluate the foundational principles of general relativity and its modifications.
@@
Soon after the works of Thorne and collaborators, Deutsch and Politzer attempted to investigate quantum effects associated with CTCs. They introduced the Deutsch-Politzer (DP) space—an abstract construct created through surgical manipulation on a two-dimensional Lorentzian spacetime, forming a cylinder-shaped ‘handle’ that bridges two spacelike regions [20, 21, 23, 22]. This process is visually depicted in Fig.  1, where Paths 3 and 4 represent CTCs. As an object moves along a CTC, proper time continues to cumulate (i.e., the time count registered by a clock carried along the path keeps increasing), but the object eventually returns to its original spacetime point. Whereas geometry pertains to local properties of spacetime, topology pertains to global ones; in crafting a ‘handle’, the DP surgery deliberately alters the topology of spacetime, allowing for CTCs to emerge.
@@
In a series of recent papers [26, 27, 28], one of the authors extended Buchdahl’s work in pure $\mathcal{R}^{2}$ gravity [29], culminating in a static spherisymmetric vacuum solution. Notably, despite the absence of exotic matter, the higher-derivative terms in the theory formally violate the WEC, under certain conditions. This violation enables the construction of a two-way traversable wormhole, which we named the Morris-Thorne-Buchdahl (MTB) wormhole [30]. Importantly, the analytical solution has allowed us to construct a ($\zeta-$)Kruskal-Szekeres (KS) diagram, revealing a non-trivial topology. The $(\zeta-$)KS coordinates enable for an identification procedure that results in the wormhole throat bridging two asymptotically flat spacetime sheets with opposing flows of proper time, as we recently proposed in [31]. Physical processes occurring in the two spacetime sheets across the throat evolve in opposite time directions. In essence, the MTB wormhole throat serves as a “gateway to another time-reversed world”. More generally, we have found that this property is a universal feature for Brans-Dicke and scalar-tensor gravities [32]. It is important to note that embedding diagrams, a standard tool to visualize wormholes, areinadequate in exposing this feature, as we shall explain in a later section of this Letter.
@@
Figure 1: The Deutsch-Politzer surgery. In the left panel, a 2-dimensional Minkowski spacetime is incised along the two dotted segments. The banks of the cuts are then glued together to form a ‘handle’, as shown in the right panel. Paths 1 and 2 represent regular trajectories in Minkowski spacetime. Path 3 is self-crossing, resulting in a portion of it forming a closed timelike curve (CTC). Path 4 depicts an infinite loop.
@@
Among a variety of solutions [27, 32], the asymptotically flat Buchdahl-inspired solution to Eq. (1) expressed in the Kruskal-Szekeres (KS) coordinates is given by [26]
@@
which exceeds $r_{\text{s}}$ when $\tilde{k}\in(-1,0)$. In this scenario, the spacetime forms a two-way traversable wormhole, referred to as the Morris-Thorne-Buchdahl (MTB) wormhole, with its throat located at $r_{*}$ given in (4), to be briefly described below. An embedding diagram is shown in Fig.  2.
@@
Identification of antipodal points and emergence of CTCs—Antipodal points play a crucial role in the $\zeta-$KS diagram. Specifically, the two thick red lines on the diagram, in Fig.  2, represent a wormhole throat, and the identification of antipodal points $(T,X)$ and $(-T,-X)$ on these lines signifies a single spacetime event occurring on the throat. It is important to note that this identification is limited strictly to points on the throat. Points that stay off the throat do not enjoy this privilege; they correspond to independent events occurring on separate spacetime sheets. (Note: Our procedure contrasts with historical maneuvers known as “elliptic/antipodal identification” [38, 36, 37, 33, 34, 35, 39]. In particular, when dealing with the regular KS diagram of a Schwarzschild spacetime, Rindler [33] assigned each antipodal pair to the same event unrestrictedly, whereas Popławski [34] applied such an assignment to antipodal points that lie on the Schwarzschild horizon.)
In Fig.  2, the trajectory $A\rightarrow B\equiv C\rightarrow D\rightarrow E\equiv F\rightarrow A$ constitutes a closed timelike loop. Note that Points B and C, as well as Points E and F, are antipodal on the throat and, therefore, identical. The segments between these points can freely take any form as long as they reside within the lightcones. One key feature is that while the segment $F\rightarrow A\rightarrow B$ progresses in an increasing order of the ‘time’ coordinate $t$, the other segment $C\rightarrow D\rightarrow E$ advances in a decreasing order of $t$. (Note: the proper time $\tau$ increases along $C\rightarrow D\rightarrow E$.) Essentially, the wormhole throat serves as a gateway to another time-reversed world by connecting two sheets with opposing physical time flows.
@@
The portal connecting two time-reversed sheets is ubiquitous not only in pure $\mathcal{R}^{2}$ gravity but also in Brans-Dicke and scalar-tensor theories, as they share a similar KS diagram [32]. It is intuitive to expect that the existence of time-reversed portals is a common feature among wormholes at large. We will now present three corollaries arising from the existence of CTCs in the MTB wormhole.
@@
Acknowledgments—H.K.N. thanks Tiberiu Harko, Mustapha Azreg-Aïnou, Luis A. Correa-Borbonet, and Nicholas Buchdahl. F.S.N.L. acknowledges support from the Fundaşão para a Ciência e a Tecnologia (FCT) Scientific Employment Stimulus contract with reference CEECINST/00032/2018, and funding from the research grants UIDB/04434/2020 and UIDP/04434/2020.

### v2

# Time-reversed information flow through a wormhole in scalar–tensor gravity

@@ Abstract
This Letter aims to advance unexplored properties of a new class of Closed Timelike Curves recently discovered in scalar–tensor gravity, reported in Universe 9, 467 (2023) and Eur. Phys. J. C 83, 626 (2023). Therein, it was shown that when the Weak Energy Condition is violated, the topology of spacetime in scalar–tensor gravity is altered, enabling the formation of two-way traversable wormholes. Furthermore, each of these wormholes acts a gateway between two time-mirrored worlds, where the two asymptotically flat sheets in the Kruskal–Szekeres diagram are glued antipodally along three directions—time $t$ and the polar and azimuth angles $(\theta,\,\varphi)$ of the 2-sphere—to form a wormhole throat. This contrasts with the standard embedding diagram which typically glues the sheets only along the $\theta$ and $\varphi$ directions. Crucially, due to the ‘gluing’ along the $t$ direction, the wormhole becomes a portal connecting the two spacetime sheets with opposite physical time flows, enabling the emergence of closed timelike loops which straddle the throat. In this Letter, we shall point out that this portal mathematically permits the possibility of backward propagation of information against time. This feature is ubiquitous for wormholes in scalar–tensor theories. In addition, we formulate the Feynman sum for transition amplitudes of microscopic particles in the proximity of a wormhole throat in which we account for timelike paths that experience time reversal.
@@
Introduction—Closed timelike curves (CTCs) are mathematical constructs first emerging in the context of General Relativity via the pioneering works of van Stockum and Gödel [1, 2]. These curves represent paths through spacetime that loop back on themselves, allowing an object or observer to return to an event in its own past. Interests in CTCs were revived in the works of Thorne and Novikov in the context of the Morris–Thorne wormholes which are thought to occur in the presence of exotic matter (or in modified gravity) [3, 4, 6, 5, 7, 8]. Despite the problematic nature of CTCs, the mathematical validity sparked considerable interest among gravitational theorists [9, 16, 12, 11, 13, 14, 10, 15, 17, 18, 19]. The concept of CTCs serves as a “gedanken experiment”, compelling us to grapple with and reevaluate the foundational principles of general relativity and its modifications.
@@
Soon after the works of Thorne and collaborators, Deutsch and Politzer attempted to investigate quantum effects associated with CTCs. They introduced the Deutsch–Politzer (DP) space—an abstract construct created through surgical manipulation on a two-dimensional Lorentzian spacetime, forming a cylinder-shaped ‘handle’ that bridges two spacelike regions [20, 21, 23, 22]. This process is visually depicted in Fig.  1, where Paths 3 and 4 represent CTCs. As an object moves along a CTC, proper time continues to cumulate (i.e., the time count registered by a clock carried along the path keeps increasing), but the object eventually returns to its original spacetime point. Whereas geometry pertains to local properties of spacetime, topology pertains to global ones; in crafting a ‘handle’, the DP surgery deliberately alters the topology of spacetime, allowing for CTCs to emerge.
@@
In a series of recent papers [26, 27, 28], one of the authors extended Buchdahl’s work in pure $\mathcal{R}^{2}$ gravity [29], culminating in a static spherically symmetric vacuum solution. Notably, despite the absence of exotic matter, the higher-derivative terms in the theory formally violate the WEC, under certain conditions. This violation enables the construction of a two-way traversable wormhole, which we named the Morris–Thorne–Buchdahl (MTB) wormhole [30]. Importantly, the analytical solution has allowed us to construct a ($\zeta-$)Kruskal–Szekeres (KS) diagram, revealing a non-trivial topology. The $(\zeta-$)KS coordinates enable for an identification procedure that results in the wormhole throat bridging two asymptotically flat spacetime sheets with opposing flows of proper time, as we recently proposed in [31]. Physical processes occurring in the two spacetime sheets across the throat evolve in opposite time directions. In essence, the MTB wormhole throat serves as a “gateway to another time-reversed world”. More generally, we have found that this property is a universal feature for Brans–Dicke and scalar–tensor gravities [32]. It is important to note that embedding diagrams, a standard tool to visualize wormholes, areinadequate in exposing this feature, as we shall explain in a later section of this Letter.
@@
Figure 1: The Deutsch–Politzer surgery. In the left panel, a 2-dimensional Minkowski spacetime is incised along the two dotted segments. The banks of the cuts are then glued together to form a ‘handle’, as shown in the right panel. Paths 1 and 2 represent regular trajectories in Minkowski spacetime. Path 3 is self-crossing, resulting in a portion of it forming a closed timelike curve (CTC). Path 4 depicts an infinite loop.
@@
Among a variety of solutions [27, 32], the asymptotically flat Buchdahl-inspired solution to Eq. (1) expressed in the KS coordinates is given by [26]
@@
which exceeds $r_{\text{s}}$ when $\tilde{k}\in(-1,0)$. In this scenario, the spacetime forms a two-way traversable wormhole, referred to as the MTB wormhole, with its throat located at $r_{*}$ given in (4), to be briefly described below. An embedding diagram is shown in Fig.  2.
@@
Identification of antipodal points and emergence of CTCs—Antipodal points play a crucial role in the $\zeta-$KS diagram. Specifically, the two thick red lines on the diagram, in Fig.  2, represent a wormhole throat, and the identification of antipodal points $(T,X)$ and $(-T,-X)$ on these lines signifies a single spacetime event occurring on the throat. It is important to note that this identification is limited strictly to points on the throat. Points that stay off the throat do not enjoy this privilege; they correspond to independent events occurring on separate spacetime sheets. (Note: Our procedure contrasts with historical maneuvers known as “elliptic/antipodal identification” [38, 36, 37, 33, 35, 34, 40, 39]. In particular, when dealing with the regular KS diagram of a Schwarzschild spacetime, Rindler [33] assigned each antipodal pair to the same event unrestrictedly, whereas Popławski and other authors [38, 36, 37, 35, 40, 39] applied such an assignment to antipodal points that lie on the Schwarzschild horizon.)
In Fig.  2, the trajectory $A\rightarrow B\equiv C\rightarrow D\rightarrow E\equiv F\rightarrow A$ constitutes a closed timelike loop. Note that Points B and C, as well as Points E and F, are antipodal on the throat and, therefore, identical. The segments between these points can freely take any form as long as they reside within the lightcones. One key feature is that while the segment $F\rightarrow A\rightarrow B$ progresses in an increasing order of the ‘time’ coordinate $t$, the other segment $C\rightarrow D\rightarrow E$ advances in a decreasing order of $t$. (Note: the proper time $\tau$ increases along $C\rightarrow D\rightarrow E$. Essentially, the wormhole throat serves as a gateway to another time-reversed world by connecting two sheets with opposing physical time flows.
@@
The portal connecting two time-reversed sheets is ubiquitous not only in pure $\mathcal{R}^{2}$ gravity but also in Brans–Dicke and scalar–tensor theories, as they share a similar KS diagram [32]. It is intuitive to expect that time-reversed portals are a common feature among wormholes in general. It is important to emphasize that the violation of the WEC plays a crucial role in forming wormholes in scalar–tensor gravity. In this situation, whereas the (local) geometry of spacetime remains Riemannian, its (global) topology has been drastically altered, which is a prerequisite for the emergence of wormholes and CTCs. This fulfills the vision of Deutsch–Politzer and Krasnikov mentioned in the Introduction [20, 21, 23, 22].
We will now present three corollaries arising from the existence of CTCs in the MTB wormhole.
@@
Interestingly, in parallel with our decoherence-based proposal that forbids the use of foreknowledge as discussed herein, a more recent article has suggested a mechanism to “erase” the memory of animate travelers, potentially avoiding time-travel paradoxes [41].
@@
As our final remark, although CTCs appear to be an inevitable feature of wormholes constructed from scalar–tensor gravity, the decoherence effect discussed herein, together with the “memory eraser” mechanism independently proposed in [41], could prevent paradoxes related to time travel of animate objects, thereby rescuing scalar–tensor gravity from inherent inconsistencies.
Acknowledgments—We wish to thank the anonymous referee for their constructive and valuable comments which helped improve the quality of our Letter.HKN thanks Tiberiu Harko, Mustapha Azreg-Aïnou, Luis A. Correa-Borbonet, Lorenzo Gavassino, and Nicholas Buchdahl. FSNL acknowledges support from the Fundaşão para a Ciência e a Tecnologia (FCT) Scientific Employment Stimulus contract with reference CEECINST/00032/2018, and funding from the research grants UIDB/04434/2020 and UIDP/04434/2020.